OpenStudy (anonymous):

Use the intercept method to graph the equation. 20x − y = −20

jimthompson5910 (jim_thompson5910):

when x = 0, what is the value of y?

OpenStudy (anonymous):

0

jimthompson5910 (jim_thompson5910):

incorrect

jimthompson5910 (jim_thompson5910):

20x - y = -20 20*0 - y = -20 ... replace x with 0 0 - y = -20 y = ???

OpenStudy (anonymous):

1 or -19

jimthompson5910 (jim_thompson5910):

0-y turns into -y

jimthompson5910 (jim_thompson5910):

so -y = -20 ----> y = ???

OpenStudy (anonymous):

-21

jimthompson5910 (jim_thompson5910):

you just change the sign of each side to get -y = -20 ----> y = 20

OpenStudy (anonymous):

Okay Jim these are the answers that I got when I worked out the problem -1,0 0,20 -0.5,10 when I graph it it goes wrong

jimthompson5910 (jim_thompson5910):

now plug in y = 0 and solve for x

OpenStudy (anonymous):

That's when I get -1,0

jimthompson5910 (jim_thompson5910):

not correct

OpenStudy (anonymous):

So I divide 20 by -20 and get -1 and since I put in 0 for y I am confused

jimthompson5910 (jim_thompson5910):

20x - y = -20 20x - 0 = -20 20x = -20 x = -1 oh sry I'm thinking of something way different

jimthompson5910 (jim_thompson5910):

so yeah you have the two points (0,20) and (-1,0)

jimthompson5910 (jim_thompson5910):

plot the two points, then draw a straight line through them to graph 20x - y = -20

OpenStudy (anonymous):

Jim are you still there?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Okay so now they did this 6x+7y=14

jimthompson5910 (jim_thompson5910):

you can use the same method to find the x and y intercepts

jimthompson5910 (jim_thompson5910):

what do you get?

OpenStudy (anonymous):

14/6 ,0 and 0,2

jimthompson5910 (jim_thompson5910):

use parenthesis when referring to ordered pairs or points

jimthompson5910 (jim_thompson5910):

also, 14/6 reduces to 7/3

jimthompson5910 (jim_thompson5910):

so we have these two points: (7/3, 0) and (0,2)

OpenStudy (anonymous):

Okay I forgot to reduce and I need to put the ordered pairs in parentheses got it